# Incremental Composition of Learned Control Barrier Functions in Unknown Environments

### v1

@@ Abstract
We consider the problem of safely exploring a static and unknown environment while learning valid control barrier functions (CBFs) from sensor data. Existing works either assume known environments, target specific dynamics models, or use a-priori valid CBFs, and are thus limited in their safety guarantees for general systems during exploration. We present a method for safely exploring the unknown environment by incrementally composing a global CBF from locally-learned CBFs. The challenge here is that local CBFs may not have well-defined end behavior outside their training domain, i.e. local CBFs may be positive (indicating safety) in regions where no training data is available. We show that well-defined end behavior can be obtained when local CBFs are parameterized by compactly-supported radial basis functions. For learning local CBFs, we collect sensor data, e.g. LiDAR capturing obstacles in the environment, and augment it with simulated data from a safe oracle controller. Our work complements recent efforts to learn CBFs from safe demonstrations—where learned safe sets are limited to their training domains—by demonstrating how to grow the safe set over time as more data becomes available. We evaluate our approach on two simulated systems, where our method successfully explores an unknown environment while maintaining safety throughout the entire execution.
@@ I Introduction
A key characteristic of a truly autonomous agent is the ability to safely explore an unknown environment by iteratively gathering data, in a safe manner, about its surroundings, and then replanning based on its new belief. Consider, as an example, a mobile search and rescue robot placed in an abandoned warehouse. Upon entering its environment, the robot has only an incomplete understanding of the obstacles in this previously unseen warehouse, and hence must explore the environment, using its onboard sensors to gather information about the surrounding obstacles. If this exploration is not done in a safe manner, the robot may accidentally enter situations where it becomes damaged and can no longer complete its mission. For real-world mission-critical deployments, handling this inherent tension between safety and liveliness in a mathematically rigorous way is of paramount importance.
@@
In this work, we attack the safe exploration problem by incrementally composing locally-learned control barrier functions (CBFs). In particular, we extend prior work in offline CBF learning [1, 2, 3, 4] to the online setting by incrementally learning local CBFs that individually ensure that the system acts safely with respect its immediate surroundings. Learning local CBFs has considerable computational advantages, especially for long-running agents, since the complexity of each learning problem scales only with the amount of new data collected, and not with cumulative data. We compose these locally-valid CBFs into a single valid global CBF via a pointwise maximum, using recent progress on non-smooth CBFs [5]. In order to ensure that our CBFs can indeed be meaningfully composed in such a manner, we utilize compactly supported basis functions [6] which ensure negative behavior off the support of the training data. Together, these ingredients allow us to incrementally expand the number of safe states that an autonomous system is able to operate in, as more data becomes available, while provably remaining safe throughout the exploration process by inheriting the rigorous safety guarantees of CBFs.
Our approach improves on existing work in the area of safe exploration with CBFs [7, 8] by leveraging a learning framework to produce dynamics-aware CBFs, rather than CBFs which only encode obstacle information, a limitation that can lead to safety-filter infeasibility or the requirement of unbounded actuation to maintain safety.
@@
Contributions: (1) We learn local CBFs that are globally well-defined by leveraging compactly supported basis functions to obtain negative end-behavior [6]. (2) We provide conditions on the basis function parameters and CBF-synthesis QP to certify that valid local CBFs compose, via pointwise maximum, to a valid, nonsmooth global CBF. (3) We provide simulations that illustrate the correctness and applicability of our method.
@@
Related Work: For a known dynamical system and environment, there is a rich body of literature on safe control synthesis via both Hamilton-Jacobi (HJ) reachability [9] and control barrier functions (CBF) [10], which are closely related [11]. Furthermore, generalizations of both HJ reachability and CBFs that handle complex system types—including multi-agent and hybrid systems—through the framework of composition have been recently explored [5, 12, 13, 14].
@@
In this work, we primarily focus on CBFs, motivated by their practical use in designing safe control laws via convex programming [10]. However, directly applying CBFs to the problem of safe exploration is difficult for a multitude of reasons. First, even when both the system and environment are known beforehand, deriving a valid CBF from first principles is a challenging problem. Recent work has focused instead on using safe expert-demonstration trajectories to learn a CBF directly from data [1, 3, 4]. While learning-based approaches significantly expand the applicability of CBFs, we cannot naïvely apply these techniques to safe exploration since we do not have a full description of the environment, nor a complete set of safe demonstration trajectories that covers the entire safe space.
@@
Recent works have addressed the issue of safe exploration with CBFs for various special cases. The work [15] considers online tuning of parameterized CBFs specialized to single-integrator, multi-agent systems. In [16], the authors show how to ensure simultaneous safety and exploration when learning a general stochastic, discrete-time system, assuming that a ground truth CBF is available a-priori. Furthermore, the works [17, 8, 7, 18, 19, 20] consider the use of sensor data (particularly LiDAR data) to obtain CBFs online. However, these works either consider specialized dynamics models [18, 17], require retraining on the entire demonstration dataset whenever new measurements are gathered [8], or restrict CBFs to e.g. elliptical sub-level sets or signed distance functions [7, 20, 19]. The latter restriction is not guaranteed to satisfy dynamic feasibility—especially under input constraints or underactuation—and hence cannot guarantee safety in general; we illustrate this issue with a Dubins car example in Section IV.
@@ II Background
Figure 1: A valid global CBF is obtained as the pointwise maximum of locally-valid CBFs. A necessary condition for this construction is that each local CBF have negative end-behavior.
@@ II-A Learning CBFs from Data
Constructing valid CBFs is generally challenging and has motivated learning-based approaches. We briefly review our prior work in [1, 2] on learning CBFs from demonstrations, which serves as a basis for this paper. Here, an expert dataset $\mathcal{D}:=\{(x_{i},u_{i})\}_{i=1}^{N}$ is given where $N$ state-action pairs demonstrate safe system behavior. From the dataset $\mathcal{D}$, a CBF of the form $h(x):=\sum_{j=1}^{J}\theta_{j}\phi_{j}(x)=\theta^{T}\phi(x)$ is learned by solving a quadratic program (QP) where $\phi(x):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{J}$ are $J$ pre-defined basis functions and $\theta\in\mathbb{R}^{J}$ are the decision variables. This QP is formed by constructing three distinct datasets using information from $\mathcal{D}$. The buffer dataset :=$\mathcal{B}:=\{(x_{i},u_{i})\}_{i=1}^{N_{b}}$ contains the boundary points of the dataset $\mathcal{D}$, obtained via boundary point detection algorithms (see [2] for details), on which we enforce the dynamics constraint (1d). The safe dataset $\mathcal{S}:=\mathcal{D}\setminus\mathcal{B}=\{(x_{i},u_{i})\}_{i=1}^{N_{s}}$ contains the remaining points in $\mathcal{D}$, on which we enforce the positivity and dynamics constraints (1a) and (1d), respectively. The unsafe dataset $\mathcal{U}:=\{x_{i}\}_{i=1}^{N_{u}}$ contains points that surround $\mathcal{D}$, e.g. obtained by sampling or again applying boundary point detection, on which we enforce the negativity constraint (1b).
@@
Given the datasets $\mathcal{S}$, $\mathcal{B}$, and $\mathcal{U}$, the hyperparameters $\gamma_{\text{safe}},\gamma_{\text{unsafe}},\gamma_{\text{dyn}}>0,b\geq 0$, a locally Lipschitz continuous extended class-$\mathcal{K}$ function $\alpha:\mathbb{R}\to\mathbb{R}$, and a basis function $\phi:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{J}$ with bounded Lipschitz constant, we learn a candidate CBF $h(x):=\phi(x)^{T}\theta-b$ by solving the convex quadratic program:
@@
$\mathrm{s.t.}\quad$ $\phi(x_{i})^{T}\theta-b\geq\gamma_{\text{safe}}$ $\mathllap{\forall x_{i}\in\mathcal{S}\quad\quad}$ (3b)
$\phi(x_{i})^{T}\theta-b\leq-\gamma_{\text{unsafe}}$ $\mathllap{\forall x_{i}\in\mathcal{U}\quad\quad}$ (3c)
@@
We refer to (3) as the Learning-CBF QP, and remark that the candidate CBF obtained by solving the Learning-CBF QP is valid when conditions on the density of the datapoints in $\mathcal{S}$, $\mathcal{B}$, and $\mathcal{U}$ and the Lipschitz constants of $h$, $f$, and $g$ are satisfied (we refer the reader to [1] for details). Specifically, we will use the Learning-CBF QP to learn local CBFs $h_{i}$ that are valid on a domain $D_{i}$. For the experiments in [1], the basis functions $\phi(x)$ are chosen to be random Fourier features [21]. Unfortunately, the oscillatory end-behavior of Random Fourier features prevents composition of multiple learned CBFs via the $\max$-operator (cf. Figure 1), since the locally learned CBFs $h_{i}$ may oscillate to positive outside of their domain $D_{i}$. Obtaining negative end-behavior motivates our use of compactly supported RBFs.
@@ II-B Compactly Supported Radial Basis Functions
Radial basis functions (RBFs) are scalar functions that exhibit radial symmetry about a center $z\in\mathbb{R}^{d_{x}}$, i.e. ${\phi}(x,z):=\bar{\phi}(\|x-z\|_{2})$ for a scalar function $\bar{\phi}:\mathbb{R}\to\mathbb{R}$. For a set of $J$ centers $Z:=(z_{1},\ldots,z_{J})$, we will combine RBFs linearly as $h(x):=\theta^{T}\phi(x,Z)-b$, where $\phi(x,Z):=({\phi}(x,z_{1}),\ldots,{\phi}(x,z_{J}))$. We consider a specific family of RBFs, referred to as compactly supported radial basis functions (CS-RBFs). CS-RBFs are zero outside of a pre-specified radius around $z$, which ensures that $h$ can be guaranteed to be negative outside of its region of validity $D$ for $b>0$. When using CS-RBFs to learn control barrier functions, we require continuous differentiability, and so we use Wendland’s functions, which are compactly supported, polynomial basis functions that are minimal in their polynomial degree [6]. Some examples of Wendland’s functions are $\bar{\phi}(r):=\max(0,1-r)^{4}(1+4r)/20$ for $d_{x}\leq 3$, and $\bar{\phi}(r):=\max(0,1-r)^{8}(3+24r+63r^{2})/5040$ for $d_{x}\leq 7$.
@@ III Safe Exploration via Incremental Composition of CBFs
Our goal is to safely guide the system to collect new measurements $M(x(t))$ in order to compose a CBF $H(x)$ that defines a safe set $S:=\{x\in\mathbb{R}^{d_{x}}:H(x)\geq 0\}$ which covers as much of $C$ as possible, while respecting $x(t)\in C$ for all times. We solve this safe exploration problem by combining local CBFs as follows. At time $t=0$, we take a measurement $M(x(0))$ and, using the Learning-CBF QP (cf. Definition 2), learn an initial CBF $h_{0}$ which is valid on some set $D_{0}\subseteq\mathbb{R}^{d_{x}}$, and renders a subset $S_{0}:=\{x\in D_{0}:h_{0}(x)\geq 0\}\subseteq C$ forward invariant. [^3] Here, $D_{0}$ is the set on which the constraints in Definition 2 are satisfied. In addition, we require that $S_{0}$ agrees with the measurement, i.e. $\text{proj}_{q}(S_{0})\subseteq M(x_{0})$. Let $k$ denote the number of measurements taken so far (starting at $k=1$). We proceed iteratively as follows:
@@
    Take a new measurement $M_{k}=M(x(t))$, and learn a CBF $h_{k}$ which is valid on some set $D_{k}\subseteq\mathbb{R}^{d_{x}}$, renders a subset $S_{k}:=\{x\in D_{k}:h_{k}(x)\geq 0\}$ forward invariant, and satisfies $\text{proj}_{q}(S_{k})\subseteq M_{k}$.
@@
In order to approach the boundary of the current invariant set $\cup_{i\in\{0,\dots,k-1\}}S_{i}$, one can choose a point $x^{-}$ such that $H_{k}(x^{-})<0$, and construct a model predictive controller (MPC) that steers $x(t)$ towards $x^{-}$. The control signal $u_{r}(t)$ generated by this MPC can then be filtered by the CBF-QP in Eq. 2 to obtain a signal $u(t)$ that safely approaches the boundary of $\cup_{i\in\{0,\dots,k-1\}}S_{i}$.
Our procedure works by learning CBFs in an incremental manner when new measurements $M_{k}$ become available. This yields an efficient algorithm where one does not need to recompute CBFs for regions of the state space where no new information about the set $C$ is obtained. However, this incremental CBF construction raises two important questions. The first question, which we address in Section III-A, is how one obtains the necessary expert dataset in Step 3) in order to apply the Learning-CBF QP while satisfying $\text{proj}_{q}(S_{k})\subseteq M_{k}$. The second question, addressed in Section III-B, is what conditions are needed to ensure that the $\max$ composition in Step 1) yields a valid global CBF.

### v2

@@ Abstract
We consider the problem of safely exploring a static and unknown environment while learning valid control barrier functions (CBFs) from sensor data. Existing works either assume known environments, target specific dynamics models, or use a-priori valid CBFs, and thus provide limited safety guarantees for general control-affine systems during exploration. We present a method for safely exploring by incrementally composing a global CBF from local CBFs. The challenge here is that local CBFs may not have well-defined end behavior outside their training domain, i.e. local CBFs may be positive (indicating safety) in regions where no training data is available. We show that well-defined end behavior can be obtained when local CBFs are parameterized by compactly-supported radial basis functions. For learning local CBFs, we collect sensor data, e.g. LiDAR capturing obstacles in the environment, and augment it with simulated data from a safe oracle controller. Our work complements recent efforts to learn CBFs from safe demonstrations, where learned safe sets are limited to their training domains, by demonstrating how to grow the safe set over time as more data becomes available. We evaluate our approach on two simulated systems, where our method successfully explores an unknown environment while maintaining safety throughout the entire execution.
@@ I Introduction
A key characteristic of a truly autonomous agent is the ability to safely explore an unknown environment by iteratively gathering data about its surroundings, in a safe manner, and then replanning based on its new belief. Consider the example of a mobile search and rescue robot placed in an abandoned warehouse. Upon entering its environment, the robot has only an incomplete understanding of the obstacles in this previously unseen warehouse, and hence must explore the environment, using its onboard sensors to gather information about the surrounding obstacles. If this exploration is not done in a safe manner, the robot could enter situations where it becomes damaged and can no longer complete its mission. For real-world, mission-critical deployments, handling this inherent tension between safety and liveliness in a mathematically rigorous way is of paramount importance.
@@
In this work, we attack the safe exploration problem by incrementally composing learned control barrier functions (CBFs). In particular, we extend prior work in offline CBF learning [1, 2, 3, 4] to the online setting by incrementally learning local CBFs that individually ensure that the system acts safely with respect to its immediate surroundings. Learning local CBFs has considerable computational advantages, especially for long-running agents, since the complexity of each learning problem scales only with the amount of new data collected, and not with cumulative data. We compose these local CBFs into a single valid global CBF via a pointwise maximum, using recent progress on non-smooth CBFs [5]. In order to ensure that our CBFs can be composed in such a manner, we utilize compactly supported basis functions [6] which ensure negative behavior off the support of the training data. Together, these ingredients allow us to expand the set of safe states that an autonomous system is able to operate in, as more data becomes available, while provably remaining safe throughout exploration by inheriting the rigorous safety guarantees of CBFs.
@@
Contributions: (1) We learn local CBFs $h_{i}$ such that each $h_{i}$ is valid on a domain $D_{i}\subset\mathbb{R}^{d_{x}}$ (locally-valid) and is negative on all $\mathbb{R}^{d_{x}}\setminus D_{i}$ (globally well-defined), by leveraging compactly supported basis functions to obtain negative end-behavior [6]. (2) We provide conditions on the basis function parameters and CBF-synthesis QP to certify that valid local CBFs compose, via pointwise maximum, to a valid, non-smooth global CBF. (3) We efficiently compose local safety computations obtained online via an oracle (e.g. Hamilton-Jacobi (HJ) reachability, model predictive controller (MPC)) through the use of local CBFs. (4) We provide simulations that illustrate the correctness and applicability of our method.
@@
Related Work: For a known dynamical system and environment, there is a rich body of literature on safe control synthesis via both HJ reachability [7] and control barrier functions (CBF) [8], which are closely related [9]. Most similar to our work is [10], which iteratively computes a HJ-reachability-based least-restrictive controller in contrast to our CBF-based safety filter; we compare the computation time of our method with this work in Section IV. Furthermore, generalizations of both HJ reachability and CBFs that handle complex system types—including multi-agent and hybrid systems—through the framework of composition have been recently explored [5, 11, 12, 13].
@@
In this work, we primarily focus on CBFs, motivated by their practical use in designing safe control laws via convex programming [8]. However, directly applying CBFs to the problem of safe exploration is difficult for a multitude of reasons. First, even when both the system and environment are known beforehand, deriving a valid CBF from first principles is a challenging problem. Recent work has focused instead on using safe expert-demonstration trajectories to learn a CBF directly from data [1, 3, 4]. While learning-based approaches significantly expand the applicability of CBFs, we cannot naïvely apply these techniques to safe exploration since we do not have a full description of the environment, nor a complete set of safe demonstration trajectories that covers the entire safe space. CBFs have also been applied in safe reinforcement learning, where the main challenge is that the system dynamics are unknown unlike in our work  [14]. We view our work as complementary as we expect our methods to compose well with online dynamics estimation.
@@
Recent works have addressed the issue of safe exploration with CBFs for various special cases. The work [15] considers online tuning of parameterized CBFs specialized to single-integrator, multi-agent systems. In [16], the authors show how to ensure simultaneous safety and exploration when learning a general stochastic, discrete-time system, assuming that a ground truth CBF is available a-priori. Furthermore, the works [17, 18, 19, 20, 21, 22] consider the use of sensor data (particularly LiDAR data) to obtain CBFs online. However, these works either consider specialized dynamics models [20, 17], require retraining on the entire demonstration dataset whenever new measurements are gathered [18], or restrict CBFs to e.g. elliptical sub-level sets or signed distance functions [19, 22, 21]. The latter restriction is not guaranteed to satisfy dynamic feasibility—especially under input constraints or underactuation—and hence cannot guarantee safety in general; we illustrate this issue with a Dubins car example in Section IV. Finally, we note that in the special case of linear dynamics with polytopic constraints, exact controlled-invariant sets can be computed [23].
@@ II Background
Figure 1: Scan $M(x(t_{3}))$ detects red hexagonal obstacle $O$, learns negative $h_{3}$ with no invariant set $S$, system remains in $\max$ CBF.
@@ II-A Learning CBFs from Data
Constructing valid CBFs is generally challenging and has motivated learning-based approaches. We briefly review our prior work in [1, 2] on learning CBFs from demonstrations, which serves as a basis for this paper. Here, an expert dataset $\mathcal{D}:=\{(x_{i},u_{i})\}_{i=1}^{N}$ is given where $N$ state-action pairs demonstrating safe system behavior. From the dataset $\mathcal{D}$, a CBF of the form $h(x):=\sum_{j=1}^{J}\theta_{j}\phi_{j}(x)=\theta^{T}\phi(x)$ is learned by solving a quadratic program (QP) where $\phi(x):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{J}$ are $J$ pre-defined basis functions and $\theta\in\mathbb{R}^{J}$ are the decision variables. This QP is formed by constructing three distinct datasets using information from $\mathcal{D}$. The buffer dataset $\mathcal{B}$ obtained via boundary point detection algorithms (see [2] for details), the safe dataset $\mathcal{S}:=\mathcal{D}\setminus\mathcal{B}$, and the unsafe dataset $\mathcal{U}$ obtained by sampling or again applying boundary point detection, together with detected obstacles.
@@
Given the datasets $\mathcal{S}$, $\mathcal{B}$, and $\mathcal{U}$, the hyperparameters $\gamma_{\text{safe}},\gamma_{\text{unsafe}},\gamma_{\text{dyn}}>0,b>0$, a locally Lipschitz continuous extended class-$\mathcal{K}$ function $\alpha:\mathbb{R}\to\mathbb{R}$, and a basis function $\phi:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{J}$ with bounded Lipschitz constant, we learn a candidate CBF $h(x):=\theta^{T}\phi(x)-b$ by solving the convex quadratic program:
@@
$\mathrm{s.t.}\quad$ $\theta^{T}\phi(x_{i})-b\geq\gamma_{\text{safe}}$ $\mathllap{\forall x_{i}\in\mathcal{S}\quad\quad}$ (3b)
$\theta^{T}\phi(x_{i})-b\leq-\gamma_{\text{unsafe}}$ $\mathllap{\forall x_{i}\in\mathcal{U}\quad\quad}$ (3c)
@@
We refer to (3) as the Learning-CBF QP, and remark that the candidate CBF obtained by solving (3) is valid when conditions on the density of points in $\mathcal{S}$, $\mathcal{B}$, and $\mathcal{U}$ and on the Lipschitz constants of $h$, $f$, and $g$ are satisfied [1]. Here, we use the Learning-CBF QP to learn local CBFs $h_{i}$ that are locally-valid on a domain $D_{i}$. In [1], the basis functions $\phi(x)$ are chosen to be random Fourier features (RFF) [24], which take the form $\phi(x)=\cos(Wx+b)$ for randomly sampled $W\in\mathbb{R}^{J\times d_{x}},b\in\mathbb{R}^{J}$, where $\cos(\cdot)$ is applied element-wise. Unfortunately, the oscillatory end-behavior of RFF prevents composition of learned CBFs via the $\max$-operator (cf. Fig. 1), since the local CBFs $h_{i}$ may oscillate to positive outside of their domain $D_{i}$ i.e., for a radius $r$ and point $x$, we can find $x^{\prime}$ with $\|x-x^{\prime}\|_{2}>r$ such that $\phi_{j}(x^{\prime})>0$. Obtaining negative end-behavior and thus global well-definedness motivates using compactly supported RBFs.
@@ II-B Compactly Supported Radial Basis Functions
Radial basis functions (RBFs) are scalar functions with radial symmetry about a center $z\in\mathbb{R}^{d_{x}}$, i.e. ${\phi}(x,z):=\bar{\phi}(\|x-z\|_{2})$ for a $\bar{\phi}:\mathbb{R}\to\mathbb{R}$. For a set of $J$ centers $Z:=(z_{1},\ldots,z_{J})$, we combine RBFs linearly as $h(x):=\theta^{T}\phi(x,Z)-b$, where $\phi(x,Z):=({\phi}(x,z_{1}),\ldots,{\phi}(x,z_{J}))$. We consider a specific family of RBFs, referred to as compactly supported radial basis functions (CS-RBFs). CS-RBFs are zero outside of a radius $s$ around $z$, which ensures that $h_{i}$ is negative outside of its validity region $D_{i}$ for $b>0$, and thus globally well-defined. When using CS-RBFs to learn CBFs, we require continuous differentiability, and so we use Wendland’s functions, which are compactly supported, polynomial basis functions that are minimal in their polynomial degree [6]. An example of a Wendland’s function is $\bar{\phi}(r):=\max(0,1-r)^{4}(1+4r)/20$ for $d_{x}\leq 3$.
@@ III Safe Exploration via Incremental Composition of CBFs
Our goal is to collect measurements $M(x(t))$ and build a composite CBF $H(x)$ that defines a safe set $S:=\{x\in\mathbb{R}^{d_{x}}:H(x)\geq 0\}$ that covers as much of $C$ as possible, while respecting $x(t)\in C$ for all times. We solve this safe exploration problem by combining local CBFs $h_{i}$ as follows. At time $t=0$, we take a measurement $M(x(0))$. We sample $M(x(0))$ (and $M(x(t_{k}))$ in future iterations) to generate a dataset $\mathcal{D}$, from which $\mathcal{S},\mathcal{B}$, and $\mathcal{U}$ are constructed as discussed in Section II-A; we discuss how to obtain control inputs for this dataset in Definition 3. Using the Learning-CBF QP (cf. Definition 2), we learn an initial CBF $h_{0}$ which is valid on some set $D_{0}\subseteq\mathbb{R}^{d_{x}}$, and renders a subset $S_{0}:=\{x\in D_{0}:h_{0}(x)\geq 0\}\subseteq C$ forward invariant. [^3] Here, $D_{0}$ is the set on which the constraints in Definition 2 are satisfied. In addition, we require that $S_{0}$ agrees with the measurement, i.e. $\text{proj}_{q}(S_{0})\subseteq M(x_{0})$. We then repeat this process. Let $k$ denote the number of measurements taken so far (starting at $k=1$), and let $t_{k}$ denote the time the $k$-th measurement is taken at. We summarize the procedure as follows:
@@
    Take a new measurement $M_{k}=M(x(t_{k}))$, and learn a CBF $h_{k}$ which is valid on some set $D_{k}\subseteq\mathbb{R}^{d_{x}}$, renders a subset $S_{k}:=\{x\in D_{k}:h_{k}(x)\geq 0\}$ forward invariant, and satisfies $\text{proj}_{q}(S_{k})\subseteq M_{k}$.
@@
The key insight of our procedure is that by ensuring that the learned invariant set $S_{k}$ satisfies $\text{proj}_{q}(S_{k})\subseteq M_{k}=M(x(t_{k}))$, we keep the system safe for all time.
@@
In order to approach the boundary of the current invariant set $\cup_{i\in\{0,\dots,k-1\}}S_{i}$, one can choose a point $x^{-}$ such that $H_{k}(x^{-})<0$, and construct an MPC that steers $x(t)$ towards $x^{-}$. The control signal $u_{r}(t)$ generated by this MPC can then be filtered by the CBF-QP in Eq. 2 to obtain a signal $u(t)$ that safely approaches the boundary of $\cup_{i\in\{0,\dots,k-1\}}S_{i}$.
Learning CBFs in an incremental manner when new measurements $M_{k}$ become available yields an efficient algorithm where one need not recompute CBFs for regions in the state space where no new information about the set $C$ is obtained. However, this incremental CBF construction raises two important questions. The first, which we address in Section III-A, is how one obtains the necessary expert dataset in Step 3) in order to apply the Learning-CBF QP while satisfying $\text{proj}_{q}(S_{k})\subseteq M_{k}$. The second, addressed in Section III-B, is what conditions are needed to ensure that the $\max$ composition in Step 1) yields a valid global CBF.